# Unsupervised anomaly detection in MeV ultrafast electron diffraction

### v1

@@ Abstract
Abstract
I Introduction
@@
MeV ultrafast electron diffraction (MUED) is a pump-probe characterization technique for studying ultrafast processes in materials. The use of relativistic beams leads to decreased space-charge effects compared to typical ultrafast electron diffraction experiments employing energies in the keV range [1, 2]. Compared to other ultrafast probes such as X-ray free electron lasers, MUED has a higher scattering cross section with material samples and allows access to higher order reflections in the diffraction patterns due to the short electron wavelengths. However, this is a relatively young technology and several factors contribute to making it challenging to utilize, such as beam instabilities which can lower the effective spatial and temporal resolution.
@@
In the past years, machine learning (ML) approaches to materials and characterization techniques have provided a new path towards unlocking new physics by improving existing probes and increasing the user’s ability to interpret data. In particular, ML methods can be employed to control characterization probes in near-real time, acting as virtual diagnostics, or ML can be deployed to extract features and effectively denoise acquired data. In this later case, convolutional neural network architectures such as autoencoder models are an attractive and more powerful .. ML has yet to be applied to the MUED technique, where it can certainly enable advances that can further our understanding of ultrafast material processes in a variety of systems.
@@
This study focus in the construction of an unsupervised anomaly detection methodology to detect faulty images in MUED. We believe that unsupervised techniques are the best choice for our purposes because the data used to train the detector does not need to be manually labeled, and instead, the machine is intended to detect by itself the anomalies in the dataset, which liberates the user of tedious, time-consuming initial image examination. The structure must, additionally, provide the user with some measure of uncertainty in the detection, so the user can take decisions based on this measure. If most images have a low uncertainty in the decision, the user may decide to manually examine a few images not confidently detected by the algorithm or directly throw them away to eliminate the associated risk of using them.
@@
Non supervised methods can be roughly grouped in density-based, bagging, deep learning or kernel methods. The residual methods can use some of the above-mentioned techniques. Density based methods are those that estimate a compact observation model and from it a likelihood of these observations can be inferred. Then, the anomalous nature of a sample can be judged from this measure. In [3] histograms are used for this purpose and in [4] dissimilarity measures between samples are computed to determine whether a sample is an outlier or not. The k-nearest neighborhood strategy is used in [5, 6] to provide an approximate measure of likelihood. The bagging models use an ensemble of detectors constructed with bagging or data subsampling [7, 8].
@@
In the field of deep learning we find methods that are also non-supervised and that are intended to process the data to reduce its dimensionality in what is called a bottleneck while preserving its distribution properties. These approaches are convenient when the dimension of the data is very high, as it is the case of images. The autoencoder model [9] is used for this purpose in [10, 11]. Generative adversarial networks [12] are applied for anomaly detection in [13].
@@
Our approach combines the use of an autoencoder and a residual method. Indeed, the method encodes and reconstructs the images with the help of a convolutional autoencoder (CAE) [21] and then, the reconstruction error is used as a residual to estimate the posterior probability of that the image is normal conditional to the observed error. The error is modeled in a probabilistic way using a convex combination of distributions that automatically capture the conditional distribution of the error under the hypothesis of normal or fault.
II Experimental Methods
The MUED instrument is located in the Accelerator Test Facility at Brookhaven National Laboratory. A schematic representation of the experimental setup is presented in Fig. 1. The femtosecond electron beams are generated using a frequency-tripled Ti:Sapphire laser that illuminates a copper photocathode generating a high brightness beam. The electrons are then accelerated and compressed in a 1.6-cell RF cavity achieving energies up to 5 MeV. Current parameters of the electron beam source optimized for stability are presented in Table 1. The sample chamber is located downstream from the source with a motorized holder for up to nine samples with cryogenic cooling capabilities and a window to allow laser pumping of the material. Next to the chamber a RF deflecting cavity is located and 4 m downstream the detector system is placed to collect the diffraction patterns. The detector consists of a phosphor screen followed by a copper mirror (with a hole for non-diffracted electrons to pass) and a CCD Andor camera of 512 pixels x 512 pixels with a large aperture lens. Suitable material systems for MUED require careful preparation with typical lateral sizes of 100 - 300 $\mu$m and roughly ¡ 100 nm thickness to assure electron transparency. Laser fluency is adjusted to avoid radiation-induced damage of the probed material.
@@ II Experimental Methods
Number of electrons per sec per $\mu$m2 88 - 880
In the present study, electron diffraction patterns were measured for a thin single crystal Ta2NiSe5 sample. No pump laser was employed and single shot diffraction patterns were captured. Figure 2(a) shows a typical diffraction pattern measured for Ta2NiSe5 consisting of a grayscale image of 512 x 512 pixels. Figures 2(b)-(d) present different types of anomalies due to instabilities in the electron beam. A typical dataset will contain between 70 - 90 single shot diffraction patterns, including such anomalies that will affect the accuracy of the experiment. For this work, several datasets were captured during multiple days of operation of the MUED experiment employing the Ta2NiSe5 sample.
@@
III Computational Methods
III.1 Data pre-processing procedure
@@
Each input pattern was normalized and divided in overlapping tiles of 80 x 80 pixels. Tiles containing only background were filtered out of the dataset. Tiles were defined as background if they contained white noise, and to evaluate this we applied the concept of spectral inverse participation ratio ($sIPR$). Given an image, we can classify it as white noise if it has a flat spatial spectral density. For simplicity, let’s evaluate the frequency distribution in one dimension with $f_{i}(x)$ the contribution to the $i$-th frequency. If this distribution is normalized, we define the $sIPR$ as:
$sIPR=\sum_{i=1}^{N}f_{i}^{2}(x)$ (1)
with $N$ the total number of discreet frequencies present in the image. If the distribution is consistent with white noise then $f_{i}(x)$ is equal for all frequencies and it follows that:
$sIPR_{white-noise}=\frac{1}{N}$ (2)
which is also the lowest possible value of the $sIPR$. Therefore, if a tile is found to have $sIPR>sIPR_{white-noise}$ it is included in the dataset. A slightly more restrictive threshold was adopted to limit those tiles with low signal-to-noise ratio with $sIPR>\frac{1.009}{N}$ being the condition for tiles to be included in the dataset.
This pre-processing procedure also served as a data augmentation technique and to decrease the computational cost of the calculations.
III.2 Convolutional autoencoder for reconstruction
@@
A CAE has been used as the main element of the system. A schematic of the CAE model used for reconstruction of the diffraction patterns can be found in Fig. 3 and the full structure and parameters are presented in Table 2. The parameters of the CAE have been cross-validated with the use of a set of normal images.
@@
This output is then passed through a nonlinear transformation with two steps, a nonlinear activation and a MaxPooling [23, 24] to produce output channel ${\bf H}^{\ell}_{j}$, that will be summarized detailed further.
The first convolution layer has 64 output channels and the size of each channel is $80\times 80$. The convolution kernel is $3\times 3$ and a zero padding of one pixel is applied to each edge of the resulting image. Therefore, the convolution output has a width and height of $80-3+1=78$. Due to the zero padding, the dimension is increased to $80$. After that, a MaxPooling is applied with a window of dimensions $4\times 4$. The MaxPooling consists of selecting the pixel with the maximum value inside a window of pixels, for all disjoint windows of the image. Therefore, the dimension of the resulting image is reduced by a factor 4 in each dimension. This results in a final dimension of $20\times 20$ pixels. After this, a Rectified Linear Unit (ReLU) activation [25] is applied. The ReLU activation is the simplest nonlinear activation and it can be written as
@@
In this operation, instead of a convolution, a Hadamard matrix product between the kernel and the image is applied. The dimension of the resulting image is equal to the product of the dimension of the input image times the dimension of the kernel. The first transpose convolution has a kernel of dimensions $5\times 5$, and since the input has a single pixel, the resulting images have dimensions $5\times 5$. The dimensionality is further increased with two more transpose convolutions up to ×$\times 80$ and 64 channels. In order to combine all the channels, a last convolution with a kernel of dimensions $4\times 4$ and zero padding with $p=1$ results in a single channel of dimension $80\times 80$. The output layer is a 2D convolution operation with a linear activation.
The mean squared error ($MSE$) loss was used with an Adam optimizer and standard backpropagation.
III.3 Anomaly detection
@@
If a significant number of anomalies are present in a dataset, the CAE will present a high training error in these images. In preliminary experiments, we observed that the errors group in two separate clusters, and therefore the likelihood function of these errors can be modeled with a mixture of probability density functions $p(e|N)$ and $p(e|A)$ [29], which are the likelihoods of the error under the hypothesis that the image is normal ($N$) or an anomaly ($A$). If the prior probability for these anomalies $p(A)=1-p(N)$, then a posterior probability can be written as
@@
can be modeled with two Gamma distributions or two Rice distributions as conditional likelihoods. If it is assumed that a significant number of anomalies is present in a dataset, but they are not previously detected, as it is the case of this work, both the parameters of the distributions and the values of priors $p(N)$ and $p(A)$ can be easily adjusted by the use of the ME algorithm or by using the limited memory Broyden–Fletcher–Goldfarb–Shanno with box constraints (L-BFGS-B) [30] algorithm.
@@
The process is visualized in Fig. 4. The process starts with the introduction of an image, which is tiled. Each of the tiles ${\bf T}_{i}$ is reconstructed and its reconstruction error ${\bf E}_{i}$ is computed. The, the logarithm of the mean square error across all the pixels of the tiles is computed and with it, an estimation of the posterior $p(N|e)=p(N|{\bf I})$ is computed. In the system, it is assumed that both likelihoods and the priors are estimated. The details of the particular algorithm for this estimation are provided in the next section.
@@
Figure 5: Comparison between 10 original and reconstructed test tiles of an image and their corresponding reconstruction error. The graphs show the logarithm of the amplitudes and the squared error.
III.4 Computational resources
@@
IV Experiments
IV.1 Training and test of the autoencoder
The autoencoder has been trained with $6537$ tiles corresponding to 100 images. The machine is agnostic about which images are faulty. The training algorithm is the standard Adam with MMSE criterion. The training parameters are: $\mu=0.01$, batch size equal to 32, and 10 training epochs. The total training time was of 17 minutes. An example of the reconstruction for 10 arbitrary normal tiles of an image is shown in Fig. 5
During the test, each image is segmented in tiles, and passed to the autoencoder. The mean squared error is computed across all pixels of all tiles of the image.
The CAE was tested with the available 1521 images (including the training ones), of which 615 (40.4%) of them were faulty images, and 906 were normal. The total number of analyzed tiles is about $10^{5}$. The total test time was of 24 minutes, corresponding roughly to 1 second per image. This includes the image loading from the hard drive, which took most of the processing time. The logarithm of the MSE across all pixels of each tested image can be seen in Fig. 6.
Figure 6: Histogram of the logarithm of the MSE of the reconstruction error for all the available images.
IV.2 Statistical analysis of the reconstruction error

### v2

@@ Abstract
MeV ultrafast electron diffraction (MUED) is a pump-probe technique used to study the dynamic structural evolution of materials. An ultrashort laser pulse triggers structural changes, which are then probed by an ultrashort relativistic electron beam. To overcome low signal-to-noise ratios, diffraction patterns are averaged over thousands of shots. However, shot-to-shot instabilities in the electron beam can distort individual patterns, introducing uncertainty. Improving MUED accuracy requires detecting and removing these anomalous patterns from large datasets. In this work, we developed a fully unsupervised methodology for the detection of anomalous diffraction patterns. Using a convolutional autoencoder, we calculate the reconstruction mean squared error of the diffraction patterns. Based on the statistical analysis of this error, we provide the user an estimation of the probability that the pattern is normal, which also allows a posterior visual inspection of the images that are difficult to classify. This method has been trained with only 100 diffraction patterns and tested on 1521 patterns, resulting in a false positive rate between 0.2% and 0.4%, with a training time of 10 seconds per image and a test time of about 1 second per image. The proposed methodology can also be applied to other diffraction techniques in which large datasets are collected that include faulty images due to instrumental instabilities.
journal: Nucl. Instr. and Meth. in Physics Res. Section A
1 Introduction
@@
MeV ultrafast electron diffraction (MUED) is a pump-probe characterization technique for studying ultrafast processes in materials [32, 35, 33, 24]. The use of relativistic beams leads to decreased space-charge effects with respect to typical ultrafast electron diffraction experiments employing energies in the keV range. Compared to other ultrafast probes such as X-ray free electron lasers, MUED has a higher scattering cross section with material samples and allows access to higher-order reflections in the diffraction patterns due to the short electron wavelengths. However, this is a relatively young technology and several factors contribute to making it challenging to utilize, such as beam instabilities which can lower the effective spatial and temporal resolution.
In a typical MUED experiment, diffraction patterns are averaged over many shots to increase the signal-to-noise ratio. However, shot-to-shot instabilities in the electron beam can distort individual patterns leading to lower resolution in the averaged diffraction patterns. These faulty or anomalous diffraction patterns, arising from drifts in the electron beam, are averaged into the dataset; in our acquisition chain, images are streamed and handled at an effective throughput of  25 Mb/s (data rate) at 5 Hz repetition. The resulting loss in resolution can obscure subtle structural changes in the material, especially in long experiments where significant drifts are expected to arise. Detection and elimination of these anomalous diffraction patterns from the large datasets generated by MUED can pave the way into increasing the resolution of the technique to explore ultrafast material processes with exceptional accuracy.
@@
In the past years, machine learning (ML) approaches to materials and characterization techniques have provided a new path towards unlocking new physics by improving existing probes and increasing the user’s ability to interpret data. In particular, ML methods can be employed to control characterization probes in near-real time, acting as virtual diagnostics, or ML can be deployed to extract features and lead to enhanced data analysis. ML up to now has yet to be applied to the MUED systems, where it can certainly enable advances that can further our understanding of ultrafast material processes in a variety of systems.
@@
This study focuses on the construction of an unsupervised anomaly detection methodology to detect faulty images in MUED. We believe that unsupervised techniques are the best choice for our purposes because the data used to train the detector does not need to be manually labeled, and instead, the machine is intended to detect the anomalies in the dataset, which liberates the user from tedious, time-consuming initial image examination. The structure must, additionally, provide the user with some measure of uncertainty in the detection, so the user can take decisions based on this measure. If most images have a low uncertainty in the decision, the user may decide to manually examine a few images not confidently detected by the algorithm or directly throw them away to eliminate the associated risk of using them.
@@
Non supervised methods can be roughly grouped in density-based, bagging, deep learning, or kernel methods. The residual methods can use some of the above-mentioned techniques. Density-based methods are those that estimate a compact observation model and from it a likelihood of these observations can be inferred. Then, the anomalous nature of a sample can be judged from this measure. In [11], histograms are used for this purpose, and in [4], dissimilarity measures between samples are computed to determine whether a sample is an outlier or not. The k-nearest neighborhood strategy is used in [25, 2] to provide an approximate measure of likelihood. The bagging models use an ensemble of detectors constructed with bagging or data subsampling [18, 1].
@@
In the field of deep learning, we find methods that are also non-supervised and that are intended to process the data to reduce its dimensionality in what is called a bottleneck while preserving its distribution properties. These approaches are convenient when the dimension of the data is very high, as it is the case of images. The autoencoder model [15] is used for this purpose in [26, 6]. Generative adversarial networks [12] are applied for anomaly detection in [20].
@@
Our approach is based on the use of a lightweight convolutional autoencoder (CAE) [21] that, with low computational burden, is able to reconstruct the input images and denoise them conveniently. The reconstruction error is used as a residual to estimate the posterior probability of that the image is normal conditional to the observed error. The error is modeled in a probabilistic way using a convex combination of distributions that automatically capture the conditional distribution of the error under the hypothesis of normal or fault. The method provides added value with respect to a standard statistical evaluation in terms of its capability to denoise the images through the use of the CAE, together with an ability to estimate the probability that an image is faulty. Then, the user can discard the images with a high probability and only manually inspect those that have a probability close to 0.5, which are, as it is seen in the experimental Section, a very small fraction of the analyzed images.
2 Experimental Methods
[…]
@@ 2 Experimental Methods
[…]
In the present study, electron diffraction patterns were measured for a thin single crystal Ta2NiSe5 sample [3, 7]. No pump laser was employed and single shot diffraction patterns were captured. Figure 2(a) shows a typical diffraction pattern measured for Ta2NiSe5 consisting of a grayscale image of 512 x 512 pixels. Figures 2(b)-(d) present different types of anomalies due to instabilities in the electron beam. A typical dataset will contain between 70 - 90 single shot diffraction patterns, including such anomalies that will affect the accuracy of the experiment. For this work, several datasets were captured during multiple days of operation of the MUED experiment employing the Ta2NiSe5 sample.
@@
3 Computational Methods
3.1 Data pre-processing procedure
@@
Each input pattern was normalized and divided in overlapping tiles of 80 x 80 pixels. Tiles containing only background were filtered out of the dataset. Tiles can be defined as background if they contain white noise. To evaluate this, it is simpler to determine whether the image has a high probability of containing Bragg peaks by counting the number of connected pixels with an amplitude over a threshold computed as a function of the median of the tile amplitude. The used algorithm simply consists of detecting all pixels in the image with a value over $m$ times the median. If a pixel is copnnected to $n$ or more other pixels over the same threshold, the image is selected. The chosen values were $m=3$ and $n=10$, which correctly selected all tiles containing Bragg peaks in a small sample. An example of the result of image tiling can be seen in Fig. 3, where the logarithm of the pixel amplitude is shown.
Figure 3: Original normal image and the corresponding tiled image after the preprocessing. The figure represents the logarithm of thre amplitude to enhance the details of the tiling.
The size of the images is set so that they are significantly larger than a Bragg spot, but small enough not to compromise the computational burden of the methodology during training. As it will be detailed further, a choice of $80\times 80$ pixels with the chosen architecture leads to a latent space of dimension 256 (see Table 2 below) after 3 convolutions. If we increase the dimension of the input, the dimension of the latent space increases quadratically, hence the computational burden.
To crossvalidate parameters $m$ and $n$, we conducted experiments with different integer values of $m$ from 1 to 4 and different values of $n$ equal to 3, 6, 10 and 20. In the 16 experiments, we visually detected the number of wrongly selected tiles among a set of 160 tiles. The combinations $m=3$ and $n=6$ or 10 detected all tiles without error.
The tile discrimination was done during the training only, but during the test, all tiles were processed in order to be able to reconstruct full denoised images.
3.2 Convolutional autoencoder for reconstruction
@@
A convolutional autoencoder (CAE) has been used as the main element of the system. A schematic of the CAE model used for reconstruction of the diffraction patterns can be found in Fig. 4 and the full structure and parameters are presented in Table 2. The parameters of the CAE have been cross-validated with the use of a set of normal images.
@@
This output is then passed through a nonlinear transformation with two steps, a nonlinear activation and a MaxPooling [10, 19] to produce output channel ${\bf H}^{\ell}_{j}$, that will be summarized further.
The first convolution layer has 64 output channels and the size of each channel is $80\times 80$. The convolution kernel is $3\times 3$ and a zero padding of one pixel is applied to each edge of the resulting image. Therefore, the convolution output has a width and height of 803+1=7880-3+1=78. Due to the zero padding, the dimension is increased to $80$. After that, a MaxPooling is applied with a window of dimensions $4\times 4$. The MaxPooling consists of selecting the pixel with the maximum value inside a window of pixels, for all disjoint windows of the image. Therefore, the dimension of the resulting image is reduced by a factor 4 in each dimension. This results in a final dimension of $20\times 20$ pixels. After this, a Rectified Linear Unit (ReLU) activation [23] is applied. The ReLU activation is the simplest nonlinear activation and it can be written as
@@
In this operation, instead of a convolution, a Hadamard matrix product between the kernel and the image is applied. The dimension of the resulting image is equal to the product of the dimension of the input image times the dimension of the kernel. The first transpose convolution has a kernel of dimensions $5\times 5$, and since the input has a single pixel, the resulting images have dimensions $5\times 5$. The dimensionality is further increased with two more transpose convolutions up to $\times 80$ and 64 channels. In order to combine all the channels, a last convolution with a kernel of dimensions $4\times 4$ and zero padding with $p=1$ results in a single channel of dimension $80\times 80$. The output layer is a 2D convolution operation with a linear activation. The choice of the parameters of the CAE has not been fully explored, due to the limitations of the computational resources at hand, but the performance of the present architecture can be improved by such crossvalidation of parameters. The use of larger images implies a higher number of parameters, at least in the input and output layers, which translates into convolutions of a much higher computational burden. With the maxpool sizes of the present architecture, it would take more convolutional layers to arrive to a small bottleneck (128 or 256). Also, the convolutions need to capture local relationships between a Bragg peak in order to proceed to the reconstruction, and therefore, we did not consider it necessary to use larger images.
The mean squared error ($MSE$) loss was used with an Adam optimizer [17] and standard backpropagation. The justification of the use of the MSE criterion is simply based on the assumption that the error is Gaussian. If the error likelihood is Gaussian and a Gaussian prior is established for the set of network parameters, then the posterior probability of these parameters given the training data can be written as $p({\bf e}|{\mathcal{X}})\propto\exp\left(-\gamma{\bm{\Theta}}\right)\prod_{i}\exp\left(\frac{-\sigma^{-2}}{2}\|{\bf X}_{i}-{\bf{\hat{X}}_{i}}\|^{2}\right)$. The first term is the prior and the second models the likelihood, where ${\bf X}_{i},~{\bf{\hat{X}}_{i}}$ represent the input and reconstructed images, respectively. The minimization of the mean square error plus a regularization term is equivalent to maximizing the log posterior of the parameters, often called Maximum a Posteriori (MAP) criterion. Indeed, taking logarithms in both sides and negating them leads to the expression $\sum_{i}\frac{\sigma^{-2}}{2}\|{\bf X}_{i}-{\bf{\hat{X}}_{i}}\|^{2}+\gamma{\bm{\Theta}}$, which contains the sum of the squared errors plus a regularization term, that is included in the Adam optimization algorithm.
3.3 Anomaly detection
@@
If a significant number of anomalies are present in a dataset, the CAE will present a high training error in these images. In preliminary experiments, we observed that the errors group in two separate clusters, and therefore the likelihood function of these errors can be modeled with a mixture of probability density functions $p(e|N)$ and $p(e|A)$ [22], which are the likelihoods of the error under the hypothesis that the image is normal ($N$) or an anomaly ($A$). If the prior probability for these anomalies p(A)=1p(N)p(A)=1-p(N), then a posterior probability can be written as
@@
can be modeled with two Gamma distributions or two Rice distributions as conditional likelihoods. If it is assumed that a significant number of anomalies is present in a dataset, but they are not previously detected, as it is the case of this work, both the parameters of the distributions and the values of priors $p(N)$ and $p(A)$ can be easily adjusted by the use of the ME algorithm or by using the limited memory Broyden–Fletcher–Goldfarb–Shanno with box constraints (L-BFGS-B) [9] algorithm. The details of the strategy to asses the convergence of the algorithm are explained in Subsection 4.2 below.
It must be noted that in dynamic experiments, we must avoid treating real pump-induced changes as anomalies. The cleanest deployment is to train (or calibrate the error distribution) on a representative subset of nominally good frames for each condition (unpumped and pumped) or on pre-pump baseline frames, and then apply the detector within each condition/time-delay block. The model’s output probability $p(N|e)$ then identifies frames that deviate from the typical appearance for that condition.
@@
The process is visualized in Fig. 5. The process starts with the introduction of an image, which is tiled. Each of the tiles ${\bf T}_{i}$ is reconstructed and its reconstruction error ${\bf E}_{i}$ is computed. Then, the logarithm of the mean squared error across all the pixels of the tiles is computed and with it, an estimation of the posterior $p(N|e)=p(N|{\bf I})$ is computed. In the system, it is assumed that both likelihoods and the priors are estimated. The details of the particular algorithm for this estimation are provided in the next section.
@@
Figure 6: Comparison between 10 original and reconstructed test tiles of an image and their corresponding reconstruction error and the entire reconstructed image. The graphs show the logarithm of the amplitudes and the squared error. The original and reconstructed features are shown in pixel coordinates.
3.4 Computational resources
@@
4 Experiments
4.1 Training and test of the autoencoder
Training is performed using 100 training images (resulting in 5492 tiles) and 30 validation images (1520 tiles). The batch size is 32 tiles, which, given the current tiling parameters, corresponds to batches of 172 tiles. The training algorithm is the standard Adam with MMSE criterion. The training parameters are: $\mu=0.01$, and 10 training epochs. The total training time was of less than 2 minutes. The machine is agnostic about which images are faulty. The mean and standard deviation of the training images are computed, and then training and validation (and later test) images are normalized with these parameters. No further preprocessing is applied to the images.
After each epoch, we evaluate performance on the 30 validation images. We observe that the validation error does not change significantly after 10 epochs; however, a more rigorous stopping strategy should rely on an explicit criterion as early stopping when the validation error reaches a minimum within a specified tolerance, or stopping upon detection of overfitting (i.e., when the validation error begins to increase).
During the test, each image is segmented in tiles, without discarding the ones not showing Bragg peaks, and passed to the autoencoder. The mean squared error is computed across all pixels of all tiles of the image. An example of the reconstruction for 10 arbitrary normal tiles of an image is shown in Fig. 6, where the original and reconstructed features are shown in pixel coordinates.
The CAE was tested with the available 1521 images (including the training ones), of which 615 (40.4%) of them were faulty images, and 906 were normal. The total number of analyzed tiles is about $10^{5}$. The total test time was of 24 minutes, corresponding roughly to 1 second per image. This includes the image loading from the hard drive, which took most of the processing time.
4.2 Statistical analysis of the reconstruction error